OpenStudy (anonymous):

The Fiedler family has up to $130,000 to invest. They decide that they want to have at least $40,000 invested in stable bonds yielding 5.5% and that no more than $60,000 should be invested in more volatile bonds yielding 11%. How much should they invest in each type of bond to maximize income if the amount in the stable bond should not exceed the amount in the more volatile bond? What is the maximum income?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

first define your variables let x = amount of money (in dollars) invested in the stable bonds at 5.5% y = amount of money (in dollars) invested in the volatile bonds at 11%

jimthompson5910 (jim_thompson5910):

They have up to $130,000 to invest, so this means that the total they have to invest cannot be larger than 130,00 so this means total <= 130000 x+y <= 130000 since x+y is the total amount to invest (in both combined)

OpenStudy (anonymous):

stable bonds = 130.000 - 60,000 = 70,000

jimthompson5910 (jim_thompson5910):

we're also told that "they want to have at least $40,000 invested in stable bonds yielding 5.5%", so we know x >= 40000 and we're given that "no more than $60,000 should be invested in more volatile bonds yielding 11%", so we can say y <= 60000

OpenStudy (anonymous):

volatile bonds = not > than 60,000

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

with me so far?

OpenStudy (anonymous):

yes @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what you do from here is you graph each inequality on the same xy plane

jimthompson5910 (jim_thompson5910):

then make notes where the points of intersections are between each line

jimthompson5910 (jim_thompson5910):

to help define the feasible region

OpenStudy (anonymous):

um okl

jimthompson5910 (jim_thompson5910):

do you know how to do that?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

ok i did it all and i got

OpenStudy (anonymous):

$60,000 in the stable bonds and $60,000 in the volatile bonds; maximum income $9900

jimthompson5910 (jim_thompson5910):

let me check

OpenStudy (anonymous):

k

jimthompson5910 (jim_thompson5910):

what points of intersection did you get

OpenStudy (anonymous):

OMG MY BATTERY IS ABOUT DIE THANK YOU SO SO SO MUCH FOR YOUR HELP THUS FAR @jim_thompson5910

jimthompson5910 (jim_thompson5910):

hopefully you can grab a charger and an outlet

jimthompson5910 (jim_thompson5910):

but basically you use the points of intersection and you plug them into 0.055x + 0.11y to determine which yields the max return

OpenStudy (anonymous):

IM ON A PLAN SO NOO OUTLETS

OpenStudy (anonymous):

PLANE

jimthompson5910 (jim_thompson5910):

oh gotcha, then perhaps when you land is a good time to work on this

OpenStudy (anonymous):

OK YEAH GREAT IDEA ;-)

OpenStudy (anonymous):

@jim_thompson5910 i need your help with this same problem.